Describe, in terms of momentum of the molecules, how a pressure is exerted on the walls of

the cylinder.


1
Expert's answer
2021-05-18T10:53:03-0400

We know that

When a molecule collides with the wall of the container a change of momentum occurs.

The molecule exerts an equal but opposite force on the wall (Newton's third law).

The pressure exerted by the gas is due to the sum of all these collision forces....Thus the average speed of the molecules will increase

Newton's third law "F_{12}=-F_{21}"

Pressure (P)="\frac{force}{Area}"

"P=\frac{F}{A}"

Where "F=\frac{dp}{dt}"

"P=\frac{dp}{Adt}\rightarrow(1)"

Pressure applied wall of cylinder according to equation (1)

momentum change with respect to time